### v2

# Latest version here

Abstract
This paper studies how imprecision in noisy signals attenuates Bayesian updating toward the prior. This phenomenon is well-known under a normal prior and normal noise, where less precise signals yield posterior means closer to the prior mean. We show this effect extends to any symmetric, log-concave prior and any symmetric, quasi-concave location experiment, using a newly introduced precision order. Our main result is that for any such prior and any signal realization, the posterior mean under location experiment $S$ is closer to the prior mean than is the posterior mean under $S^{\prime}$, if and only if $S$ is less precise than $S^{\prime}$. We discuss applications to cognitive imprecision, prior precision, and overconfidence.
Keywords: signal extraction, reversion to the mean, underreaction, central tendency effect, cognitive imprecision
JEL Classifications: C11, D83, D84
1 Introduction
In many economic contexts, agents fail to respond optimally to fundamental variables: firms and households adjust only partially to macroeconomic changes (Sims, 2003), consumers underreact to non-salient taxes (Chetty, Looney, and Kroft, 2009), and numerical estimates are biased toward default values (Tversky and Kahneman, 1974). [^3] Outside economics, psychophysics documents a central-tendency effect: perceptual judgments are biased toward the mean of the stimulus distribution (Hollingworth, 1910; Stevens and Greenbaum, 1966).
Such attenuated behavior is often explained through imprecise information. In one version of this explanation, agents base their choices on noisy signals of the fundamentals, where noise arises from observational errors. In a second version, agents observe the fundamentals but the cognitive process that translates fundamentals to actions is noisy, reflecting cognitive constraints. The latter explanation has gained attention in the growing literature on cognitive imprecision (Gabaix, 2019; Woodford, 2020; Enke and Graeber, 2023; Enke, Graeber, Oprea, and Yang, 2024), which shows how noisy information processing can account for a wide range of economic phenomena.
These explanations share a core mechanism: noise attenuates Bayesian updating, which compresses behavior toward a default action. This mechanism is usually formalized using the normal-normal model, where the agent observes a noisy signal $S=X+\varepsilon$. Here, the state $X\in\mathbb{R}$ is normally distributed, and $\varepsilon\in\mathbb{R}$ is independent, normal noise. The agent’s posterior mean is then compressed toward the prior mean, and more so when the signal is less precise: the posterior mean lies between the signal and the prior mean, and the posterior mean is closer to the prior mean the larger the variance of $\varepsilon$ [^4] Formally, if ${X}\sim\mathcal{N}(\mu,\sigma_{X}^{2})$ and $\varepsilon\sim\mathcal{N}(0,\sigma_{\varepsilon}^{2})$, then $\mathbb{E}[{X}|S=s]=\frac{\sigma_{{X}}^{2}}{\sigma_{X}^{2}+\sigma_{\varepsilon}^{2}}s+\left(1-\frac{\sigma_{{X}}^{2}}{\sigma_{{X}}^{2}+\sigma_{\varepsilon}^{2}}\right)\mu.$ For any signal realization $s$, the larger $\sigma_{\varepsilon}^{2}$, the closer the posterior mean is to the prior mean. We refer to this effect as imprecision attenuates updating. If behavior is based on the posterior mean, then such attenuation of updating leads to attenuation of behavior. In fact, a key piece of evidence for cognitive imprecision relies on this effect: subjects reporting higher cognitive uncertainty exhibit more attenuated behavioral responses (Enke and Graeber, 2023; Enke, Graeber, Oprea, and Yang, 2024).
Despite its empirical relevance, it remained unknown whether the attenuation effect extends beyond the normal distributions, which rely on strong parametric assumptions. While normality can be justified in specific contexts—e.g., via the central limit theorem or rational inattention with a normal prior and quadratic loss—these cases are limited. A non-parametric class of signal structures that preserves the attenuation effect would offer a more robust foundation for interpreting empirical evidence as a consequence of imprecise information.
To address this gap, we prove the imprecision-attenuates-updating effect holds in all additive-noise models (location experiments) with symmetric, quasi-concave (unimodal) noise under a symmetric, log-concave prior: for any signal realization, the posterior mean moves toward the prior mean as the signal becomes less precise (Theorem 1).
To formalize this result, we introduce a notion of precision, which we call the precision order. This order requires the noise distribution of a less precise signal to be farther from zero in the sense of likelihood-ratio domination: $\tilde{S}=X+\tilde{\varepsilon}$ is less precise than $S=X+\varepsilon$ if the likelihood ratio $f_{\tilde{\varepsilon}}(x)/f_{\varepsilon}(x)$ increases as $x$ moves away from 0 (see Figure 1). Concretely, we show that scaling up a log-concave noise term produces a less precise signal under this definition. As a result, the precision order applies to log-concave location-scale experiments. [^5] Furthermore, we demonstrate that a weaker condition than log-concavity suffices, which holds for many non-log-concave noise distributions (Lemma 1).
We show that the precision order is not only sufficient but also necessary for the imprecision-attenuates-updating effect. The precision order is thus characterized by this effect. Specifically, if a signal structure $S$ is not more precise than another signal structure $S^{\prime}$, then there exists a symmetric, log-concave prior and a signal realization where $S$ induces more attenuation than $S^{\prime}$. This holds even if $S$ dominates $S^{\prime}$ in the Blackwell or Lehmann orders, which we show do not imply the precision order.
Symmetric, quasi-concave distributions include many common probability distributions beyond the normal, such as logistic, extreme value, and double-exponential distributions. Our focus on this class follows Chambers and Healy (2012), who demonstrate symmetry and quasi-concavity are necessary for the posterior mean to robustly lie between the prior mean and the signal realization. More precisely, they show that if the prior is not symmetric and quasi-concave, one can find a symmetric and quasi-concave noise density (and vice versa) under which the posterior mean does not lie between the prior mean and the signal realization.
Together with Chambers and Healy (2012), this paper shows that the attenuation effect holds under broader conditions than previously recognized. But the effect is not universal. As Chambers and Healy (2012) show, without symmetry and quasi-concavity, agents may overreact—the posterior mean can lie farther from the prior mean than the signal. Moreover, as we show, greater informativeness in the Blackwell sense does not guarantee less attenuation: for some signal realizations, more informative signals can lead to more attenuated updating.
Finally, we leverage Theorem 1 to derive implications of broader economic relevance for signal-extraction problems:
  • Prior Precision: Our main theorem implies a converse comparative-statics result regarding prior precision (Corollaries 2 and 3): increasing the precision of the prior brings the posterior mean closer to the prior mean for any signal realization.
  • Average Posterior Means: We extend our analysis to average posterior means conditional on the true state, which are relevant when agents acquire conditionally independent signals about a common state. Under symmetry and quasi-concavity, conditional on the true state, the average posterior mean lies between the state and the prior mean (Proposition 1).
  • Comparative Statics for Average Posterior Mean: Greater (over)confidence in the signal brings the average posterior mean closer to the true state (Proposition 2), whereas greater prior precision brings it closer to the prior mean (Proposition 3).
Related Literature
This paper introduces a new information order and provides comparative-statics results for signal structures with additive independent noise, commonly referred to as location experiments.
Prior work has explored orderings on location experiments related to the value of information. Boll (1955) shows that one location experiment Blackwell-dominates another if and only if the latter’s noise term can be derived by adding an independent noise term to the former’s. This condition is very restrictive, making location experiments comparable only in limited cases. In response, Lehmann (1988) introduced the less restrictive Lehmann order to rank location experiments for monotone decision problems. Our precision order is distinct from the Blackwell order, neither implying nor being implied by it. The precision order strengthens the Lehmann order.
Other comparative statics results have been established for location experiments with log-concave noise distributions, which satisfy the monotone likelihood-ratio property. Milgrom (1981) shows that the strict monotone likelihood-ratio property ensures higher signals constitute “good news,” in the sense of producing first-order stochastically dominant posteriors, for any prior. [^6] See also Chambers and Healy (2011) for a strengthening of Milgrom’s result and Heinsalu (2020) for an extension and economic application of Chambers and Healy (2011). Conversely, Dawid (1973) examines violations of the “good news”-property under heavy-tailed noise distributions, showing that if the noise is sufficiently heavy-tailed relative to the prior, extreme observations are “rejected,” yielding posteriors that revert to the prior (for further discussion, see O’Hagan and Pericchi, 2012). Together with our results, this highlights log-concave location-scale experiments are a well-behaved class of signal structures that retains key properties of the normal-normal model. Kartik, Lee, and Suen (2021) provides a comparative statics result with economic applications for agents with heterogeneous priors, termed “information validates the prior,” in which a more informative experiment brings another agent’s posterior mean closer to one’s prior mean.
2 Model
We restrict attention to signal structures, where the signal equals the one-dimensional state plus some independent noise, which are called location experiments. Formally, the random signal $S$ is a location experiment if it is the sum of the random state $X$ and random noise $\varepsilon$ that is independent of ${X}$, which we write as $\varepsilon\perp\!\!\!\perp{X}$. We assume the noise $\varepsilon$ has mean 0; that is, the signal is unbiased. If not, one could easily derive an unbiased signal by subtracting the mean of $\varepsilon$ from $S$.
Assumption 1 (Location Experiment).
$S={X}+\varepsilon$ (1)
$\varepsilon\perp\!\!\!\perp{X}$ (2)
$\mathbb{E}[\varepsilon]=0$ (3)
Further, we make the technical assumption that the state ${X}$ and noise $\varepsilon$ admit positive, continuously differentiable densities $f_{X}$ and $f_{\varepsilon}$ with finite absolute first moments. This guarantees finite conditional expectation $\mathbb{E}[{X}|S=s]$. We call $f_{X}$ the prior density and $f_{\varepsilon}$ the noise density.
Assumption 2.
The random state ${X}$ and noise $\varepsilon$ admit positive, continuously differentiable densities $f_{X},f_{\varepsilon}$ and have finite absolute first moments.
Following Chambers and Healy (2012), we assume symmetric and quasi-concave prior and noise densities. [^7] Quasi-concave densities are also called unimodal densities. They show that without these assumptions, the posterior mean does not necessarily lie between the prior mean and the signal realization. We believe that our comparative statics result is only interesting when the posterior mean is attenuated in the first place.
Assumption 3.
The prior and noise densities $f_{X}$ and $f_{\varepsilon}$ are symmetric and quasi-concave.
In particular, the prior density $f_{X}$ is symmetric around $\mathbb{E}[{X}]$ and the noise density $f_{\varepsilon}$ is symmetric around zero by (3).
Chambers and Healy (2012) show the following result, which we include for later reference.
Fact 1 (Chambers and Healy, 2012, Proposition 3).
Under Assumptions 1 to 3, for any signal realization $s$, the posterior mean $\mathbb{E}[{X}|S=s]$ lies weakly between the prior mean $\mathbb{E}[{X}]$ and the signal $s$. Formally,
$\forall s\leq\mathbb{E}[X]\colon\qquad s\leq\mathbb{E}[X|S=s]\leq\mathbb{E}[X],$
$\forall s\geq\mathbb{E}[X]\colon\qquad s\geq\mathbb{E}[X|S=s]\geq\mathbb{E}[X].$
Throughout the paper, we maintain Assumptions 1 to 3. In particular, all considered prior and noise densities are assumed to be positive, symmetric, and quasi-concave, unless otherwise specified.
3 Precision Order
One challenge to formalizing the idea that imprecision attenuates updating is to find the right order on signal structures. We show that the following new order, which we call the precision order, is characteristic of our desired result. Below, we discuss the relation between the precision order and well-known information orders like the Blackwell and Lehmann orders.
We define the precision order for the class of location experiments that satisfy Assumptions 1 to 3. We first define this order on noise terms (or, more generally, on random variables) and then extend it to the associated location experiments.
Definition 1 (Precision Order).
Let $\tilde{\varepsilon}$ and $\varepsilon$ be random variables with positive, symmetric-around-0, and quasi-concave densities $f_{\tilde{\varepsilon}}$ and $f_{\varepsilon}$, respectively. We say $\tilde{\varepsilon}$ is less precise than $\varepsilon$ if the likelihood ratio
$\frac{f_{\tilde{\varepsilon}}(x)}{f_{\varepsilon}(x)}$
is nondecreasing in $x$ for $x>0$. Further, we say the location experiment $\tilde{S}=X+\tilde{\varepsilon}$ is less precise than location experiment $S=X+\varepsilon$ if $\tilde{\varepsilon}$ is less precise than $\varepsilon$.
Note that by the symmetry of densities, the definition implies the likelihood ratio ${f_{\tilde{\varepsilon}}(x)/f_{\varepsilon}(x)}$ is weakly decreasing in $x$ for $x<0$. Thus, the precision order requires that, for positive values, $\tilde{\varepsilon}$ likelihood-ratio dominates $\varepsilon$ and, for negative values, $\tilde{\varepsilon}$ is likelihood-ratio dominated by $\varepsilon$. In other words, the less precise location experiment has a noise term further away from 0 in the sense of likelihood-ratio domination. Figure 1 gives an example of two noise distributions ranked by the precision order. The more precise distribution has a higher density at zero, which falls faster as we move away from zero.
Figure 1: The densities of $\varepsilon\sim\mathcal{N}(0,1)$ and $\tilde{\varepsilon}\sim\mathcal{N}(0,1.5^{2})$: the former distribution is more precise.
A natural question is whether the precision order holds under natural operations on the noise densities. In section 4.1, we show that when we scale up a symmetric, log-concave density by a constant $k>1$, the resulting density is less precise.
How does the precision order relate to existing information orders? The precision order is neither implied by nor implies the Blackwell order. To see why the precision order is not implied by the Blackwell order, consider the latter’s restrictive nature for location experiments. As Boll (1955) shows, two location experiments are Blackwell ordered if only if the noise term of one experiment can be obtained from the other’s noise term by adding independent noise. Cramér’s decomposition theorem further implies that a normal location experiment (i.e., $S=X+\varepsilon$, $\varepsilon\sim\mathcal{N}(\mu,\sigma^{2})$) Blackwell dominates another location experiment only if the latter is also normal (Cramér, 1936). By contrast, the precision order allows for normal location experiments to be less precise than experiments with non-normal noise. On the other hand, the precision order does not imply the Blackwell order, as illustrated by the noise term $\tilde{\varepsilon}=\varepsilon+\varepsilon^{\prime}$, that is the independent sum of a standard Cauchy $\varepsilon$ and $\varepsilon^{\prime}\sim U[-1,1]$. Computing the density of $\tilde{\varepsilon}$ shows that it is not less precise than $\varepsilon$.
[…]
4 Results
Our main result establishes that imprecision attenuates updating if and only if location experiments are ranked by the precision order, under the additional assumption of a log-concave prior. Recall that throughout the paper, all considered location experiments are assumed to satisfy Assumptions 1 to 3.
Theorem 1 (Imprecision Attenuates Updating).
Let the prior density be log-concave. Consider two location experiments $S=X+\varepsilon$ and $\tilde{S}=X+\tilde{\varepsilon}$. For every signal realization, the posterior mean under $\tilde{S}$ is closer to the prior mean than is the posterior mean under $S$, if and only if $\tilde{S}$ is less precise than $S$. Formally,
$\begin{split}&\forall s\leq\mathbb{E}[X]\colon\qquad\mathbb{E}[{X}|S=s]\leq\mathbb{E}[{X}|\tilde{S}=s]\leq\mathbb{E}[{X}]\\ &\forall s\geq\mathbb{E}[X]\colon\qquad\mathbb{E}[{X}|S=s]\geq\mathbb{E}[{X}|\tilde{S}=s]\geq\mathbb{E}[{X}]\end{split}$ (4)
if and only if $\tilde{S}$ is less precise than $S$.
An immediate implication of this result is as follows. Consider two agents who observe the same signal, such as an economic forecast, but differ in their assessment of the signal precision. The agent who assumes greater precision will form a posterior mean closer to the signal, regardless of the signal realization.
The proof, relegated to the Appendix, demonstrates necessity and sufficiency of the precision order. Necessity is the easier direction and shown by constructing a counterexample when two location experiments are not precision-ordered. Sufficiency is more involved. We can suppose without loss that the signal realization $s$ is larger than the prior mean. To the right of $s$, the more precise signal results in a likelihood-ratio dominated posterior, lowering decreases the posterior mean. To the left of the $s$, the more precise signal results in a likelihood-ratio dominant posterior, raising the posterior mean. The proof establishes that the former effect dominates, using a sequence of inequalities and bounds, leveraging symmetry, quasi-, and log-concavity assumptions.
For this as well as the following results, an analogous strict version of the result holds. We can require that the noise term is replaced by a strictly less precise noise in the sense that the likelihood ratio in Definition 1 is strictly decreasing. Then, for signal realizations distinct from the prior mean, the posterior mean is strictly closer to $s$.
Next, we show that our precision order arises endogenously in location-scale experiments with log-concave, and many non-log-concave, noise densities.
4.1 Location-Scale Experiments
We introduce to our location experiment a scale parameter $\sigma\in\mathbb{R}_{\geq 0}$ that scales the noise term, such that
$S_{\sigma}={X}+\sigma\varepsilon.$ (5)
We show comparative statics on the scale parameter $\sigma$.
Lemma 1.
If $\varepsilon$ is symmetric around 0 and $\log f_{\varepsilon}(e^{x})$ is concave, then $\forall\sigma^{\prime}>\sigma>0$, $\sigma^{\prime}\varepsilon$ is less precise than $\sigma\varepsilon$. In particular, $\log f_{\varepsilon}(e^{x})$ is concave if the density $f_{\varepsilon}(x)$ is log-concave. [^10] This can be seen easily if the log-density is differentiable. Define $\phi=\log f$ and $\psi=\exp$. Then, $\phi$ is concave and decreasing for positive values and $\psi$ is convex, increasing, and obtains positive values only, which implies $(\phi\circ\psi)^{\prime\prime}(x)=\psi^{\prime\prime}(x)(\phi^{\prime}\circ\psi)(x)+(\psi^{\prime}(x))^{2}(\phi^{\prime\prime}\circ\psi)(x)\leq 0$.
In the Appendix, we show that Lemma 1 follows from a known result. Many commonly used distributions are symmetric and log-concave, such as the normal, logistic, extreme value, and double-exponential distributions. Further, we give examples of symmetric and non-log-concave distributions for which $\log f_{\varepsilon}(e^{x})$ is nevertheless concave, such as the Student-t, Cauchy, and the “double” Pareto distribution.
Together, Theorem 1 and Lemma 1 imply the following important result. By Lemma 1, the result still holds if we weaken the assumption that the noise density $f_{\varepsilon}$ is log-concave to $\log f_{\varepsilon}(e^{x})$ being concave.
Corollary 1.
Let the prior and noise densities be log-concave. The posterior mean is weakly closer to the prior mean under a larger scale parameter, for any signal realization $s$. Formally, if $\tilde{\sigma}>\sigma>0$, then
$\forall s\leq\mathbb{E}[X]\colon\qquad\mathbb{E}[{X}|S_{\sigma}=s]\leq\mathbb{E}[{X}|S_{\tilde{\sigma}}=s]\leq\mathbb{E}[{X}],$
$\forall s\geq\mathbb{E}[X]\colon\qquad\mathbb{E}[{X}|S_{\sigma}=s]\geq\mathbb{E}[{X}|S_{\tilde{\sigma}}=s]\geq\mathbb{E}[{X}].$
The next section shows comparative statics results for changing the prior instead of the noise, exploiting a symmetry in location experiments.
4.2 Comparative Statics on the Prior
For location experiments, the posterior density is symmetric in the prior and in the noise density. This follows from the more general property of Bayesian updating that the posterior is proportional to the product of the prior and the likelihood, $p(x|s)\propto p(x,s)=p(x)p(s|x)$. In the case of location experiments, this implies the density of the posterior conditional on $S=s$ is proportional to $f_{X}(x)f_{\varepsilon}(s-x)$. Using this insight, Theorem 1 immediately implies a dual result for making the prior more precise.
Corollary 2.
Let the noise density be log-concave. The posterior mean is weakly closer to the prior mean under a more precise prior, for any signal realization $s$. Formally, if $\mathbb{E}[{X}]=\mathbb{E}[\tilde{X}]$ and $\tilde{X}$ is more precise than ${X}$, then
$\forall s\leq\mathbb{E}[X]\colon\qquad\mathbb{E}[X|X+\varepsilon=s]\leq\mathbb{E}[\tilde{X}|\tilde{X}+\varepsilon=s]\leq\mathbb{E}[X],$
$\forall s\geq\mathbb{E}[X]\colon\qquad\mathbb{E}[X|X+\varepsilon=s]\geq\mathbb{E}[\tilde{X}|\tilde{X}+\varepsilon=s]\geq\mathbb{E}[X].$
To illustrate Corollary 2, suppose that two agents observe the same signal about the state but one agent has a more precise prior (but with the same mean). Corollary 2 implies that the agent with the more precise prior has a posterior mean that is closer to the prior mean, for any signal realization.
By the same argument, an analogous dual result to Corollary 1 holds for scaling the prior instead of the noise density. If ${X}$ has density $f_{X}(x)$, then $k{X}$ has density $1/kf_{X}(x/k)$.
Corollary 3.
Let the prior and noise densities be log-concave. For any signal realization $s$, the posterior mean is weakly closer to the prior mean if we scale down the prior. Formally, if $0<\tilde{k}<k$ and we normalize the prior mean to zero, $\mathbb{E}[X]=0$, then
$\forall s\leq\mathbb{E}[X]\colon\qquad\mathbb{E}[kX|kX+\varepsilon=s]\leq\mathbb{E}[\tilde{k}X|\tilde{k}X+\varepsilon=s]\leq\mathbb{E}[X],$
$\forall s\geq\mathbb{E}[X]\colon\qquad\mathbb{E}[kX|kX+\varepsilon=s]\geq\mathbb{E}[\tilde{k}X|\tilde{k}X+\varepsilon=s]\geq\mathbb{E}[X].$
4.3 Average Posterior Means
Our previous comparative statics results hold for any signal realization and thus speaks to situations where agents observe the same signal. However, in many situations, agents observe distinct signals from the same signal structure, that is their signal realization are independent conditional on the state. Further, often we do not observe agent’s signal realizations but only the average posterior mean, such as when we observe only aggregate behavior from a population of individuals. What can be said about comparative statics with respect to the average posterior means given some true state ${X}$?
Before we prove comparative statics results, we show that the average posterior mean necessarily lies between the state ${X}$ and the prior mean $\mathbb{E}[{X}]$, extending Fact 1.
Proposition 1.
For any state ${x}$, the conditional average posterior mean $\mathbb{E}[\mathbb{E}[{X}|S]|{X}=x]$ lies weakly between the state $x$ and the prior mean $\mathbb{E}[{X}]$. Formally,
$\forall x\leq\mathbb{E}[X]\colon\qquad x\leq\mathbb{E}[\mathbb{E}[X|S]|X=x]\leq\mathbb{E}[X],$
$\forall x\geq\mathbb{E}[X]\colon\qquad x\geq\mathbb{E}[\mathbb{E}[X|S]|X=x]\geq\mathbb{E}[X].$
Figure 2: Illustrating the proof of Proposition 1.
The proof is illustrated using Figure 2. Let ${X}>\mathbb{E}[{X}]=0$. Conditional on state ${X}$, the distribution of the signal $S$ is is symmetric around ${X}$ so its expectation equals ${X}$. Taking the conditional expectation $\mathbb{E}[{X}|S]$ moves the distribution closer to $\mathbb{E}[{X}]$ as indicated by the arrows. Given our assumptions, the conditional expectation is antisymmetric and $s>0$ has a higher likelihood than $-s<0$. Thus, the overall effect on the average posterior mean is negative and $\mathbb{E}[\mathbb{E}[{X}|S]|{X}]<\mathbb{E}[S|{X}]={X}$. Further, because the density of $s$ is larger than the density of $-s$, integrating over all $s$ leads to a positive expectation, so $\mathbb{E}[{X}]<\mathbb{E}[\mathbb{E}[{X}|S]|{X}]$.
Overconfidence
First, we prove a comparative statics result for overconfidence in the signal. We consider two agents, $A$ and $B$, that face the same objective signal structure $S={X}+\varepsilon$ but update differently because they have different confidence in the signal. That is, agent $i\in\{A,B\}$ forms their conditional expectation $\mathbb{E}_{i}[{X}|S=s]$ as if $S={X}+\varepsilon_{i}$, $\mathbb{E}_{i}[X|S=s]:=\mathbb{E}[X|X+\varepsilon_{i}=s]$. Especially empirically relevant is the case overconfidence in the signal, also called overprecision, which is pervasive (Moore and Healy, 2008). We define a relative notion of overconfidence by generalizing the definition in Ortoleva and Snowberg (2015), which is based on the normal-normal model, using our Definition 1.
Definition 2.
$A$ is more confident than $B$ in the signal if $\varepsilon_{A}$ is more precise than $\varepsilon_{B}$.
Using this definition, we prove the following comparative statics result.
Proposition 2.
Let the prior density be log-concave and $A$ be more confident than $B$. Conditional on any state $x$, the average posterior mean of $A$ is weakly closer to the state than the average posterior mean of $B$. Formally,
$\forall x\leq\mathbb{E}[X]\colon\qquad x\leq\mathbb{E}[\mathbb{E}_{A}[{X}|S]|{X}=x]\leq\mathbb{E}[\mathbb{E}_{B}[{X}|S]|{X}=x],$
$\forall x\geq\mathbb{E}[X]\colon\qquad x\geq\mathbb{E}[\mathbb{E}_{A}[{X}|S]|{X}=x]\geq\mathbb{E}[\mathbb{E}_{B}[{X}|S]|{X}=x].$
Figure 3: Illustrating the proof of Proposition 2
This shows that more overconfident agents have, on average, posterior means closer to the state ${X}$ and further away from their prior mean $\mathbb{E}[{X}]$.
The proof builds on the proof of Proposition 1 and is illustrated using Figure 3. Because $A$ is more overconfident in the signal, their posterior mean is closer to the signal for any signal realization $s$ as well as $-s$. For positive ${X}$, because $s>0$ is more likely than $-s$, the overall effect on the average posterior mean is positive. Then, the result follows from Proposition 1.
Prior Precision
Second, we show comparative statics with respect to the prior precision. Given some state ${X}$, consider two agents, $A$ and $B$, with symmetric and quasi-concave priors that have the same mean, where agent $A$’s prior is more precise than $B$’s. Then, we have from the argument in Proposition 2 and the Corollary 2, we immediately obtain the following result.
Proposition 3.
Let the noise density be log-concave and $A$’s prior be less precise than $B$’s. Conditional on any state ${x}$, the average posterior mean of $A$ is weakly closer to the state than the average posterior mean of $B$. Formally,
$\forall x\leq\mathbb{E}[X]\colon\qquad x\leq\mathbb{E}[\mathbb{E}_{A}[{X}|S]|{X}=x]\leq\mathbb{E}[\mathbb{E}_{B}[{X}|S]|{X}=x],$
$\forall x\geq\mathbb{E}[X]\colon\qquad x\geq\mathbb{E}[\mathbb{E}_{A}[{X}|S]|{X}=x]\geq\mathbb{E}[\mathbb{E}_{B}[{X}|S]|{X}=x].$
5 Conclusion
In this paper, we extended the attenuation effect of imprecise information beyond the normal-normal model. By introducing a new order of precision, based on likelihood-ratio dominance, we demonstrated that imprecision attenuates expectations toward the prior mean across a broad class of distributions commonly used in economic modeling. This generalization provides a more robust theoretical foundation for interpreting empirical observations of attenuated behavior as resulting from imprecision, whether observational or cognitive.
Our findings also have broader implications for signal-extraction problems. We established comparative statics results regarding prior precision, showing that increased prior precision shifts the posterior mean toward the prior mean, for any given signal realization. Additionally, we analyzed how overconfidence and prior precision affect the average posterior mean. Perhaps surprisingly, our results show that the posterior means of overconfident agents are on average closer to the truth while those of agents with more precise priors are further away from it.